# GazeGen: Gaze-Driven User Interaction for Visual Content Generation

### v1

@@ Abstract
We present GazeGen, a user interaction system that generates visual content (images and videos) for locations indicated by the user’s eye gaze. GazeGen allows intuitive manipulation of visual content by targeting regions of interest with gaze. Using advanced techniques in object detection and generative AI, GazeGen performs gaze-controlled image adding/deleting, repositioning, and surface material changes of image objects, and converts static images into videos. Central to GazeGen is the DFT Gaze (Distilled and Fine-Tuned Gaze) agent, an ultra-lightweight model with only 281K parameters, performing accurate real-time gaze predictions tailored to individual users’ eyes on small edge devices. GazeGen is the first system to combine visual content generation with real-time gaze estimation, made possible exclusively by DFT Gaze. This real-time gaze estimation enables various visual content generation tasks, all controlled by the user’s gaze. The input for DFT Gaze is the user’s eye images, while the inputs for visual content generation are the user’s view and the predicted gaze point from DFT Gaze. To achieve efficient gaze predictions, we derive the small model from a large model (10x larger) via novel knowledge distillation and personal adaptation techniques. We integrate knowledge distillation with a masked autoencoder, developing a compact yet powerful gaze estimation model. This model is further fine-tuned with Adapters, enabling highly accurate and personalized gaze predictions with minimal user input. DFT Gaze ensures low-latency and precise gaze tracking, supporting a wide range of gaze-driven tasks in AR environments. Leveraging these precise gaze predictions, GazeGen facilitates visual content generation through diffusion processes, allowing users to intuitively manipulate visual content by targeting regions with their gaze. Additionally, it enables real-time object detection by focusing on specific regions indicated by the user’s gaze, improving responsiveness. We validate the performance of DFT Gaze on AEA and OpenEDS2020 benchmarks, demonstrating low angular gaze error and low latency on the edge device (Raspberry Pi 4). Furthermore, we describe applications of GazeGen, illustrating its versatility and effectiveness in various usage scenarios.
@@
Figure 2: Extended applications of gaze-driven interaction with GazeGen. (1) Real-Time Gaze Estimation: Continuous tracking of eye movements for precise gaze estimation. (2) Gaze-Driven Detection: Detecting and identifying objects based on where the user is looking. (3) Gaze-Driven Image Editing: Dynamic editing tasks such as Addition (adding objects based on the user’s gaze), Deletion/Replacement (removing or replacing objects based on the user’s gaze), Reposition (move objects by first gazing at the initial position, then the new position), and Material Transfer (change an object’s style or texture by first gazing at a reference object, then applying the style to the target object). (4) Gaze-Driven Video Generation: Creating and manipulating video content driven by the user’s gaze.
@@ 1 Introduction
With this robust gaze estimation foundation, GazeGen extends its capabilities to real-time object detection by leveraging gaze points to focus on specific regions of the image, retrieving object categories and bounding boxes. For visual content generation, GazeGen uses gaze as a natural command for dynamic image editing and video creation, enabling intuitive operations such as addition, deletion, repositioning, and material transfer. This comprehensive approach allows users to seamlessly manipulate visual content through their gaze, setting a new standard for accessibility and efficiency in the field.
@@
    Use of Eye Gaze for Visual Content Manipulation: We are the first to propose using eye gaze for comprehensive visual content generation and editing, such as adding, deleting, repositioning elements, material transfer, and generating videos. Additionally, GazeGen can detect and interact with objects based on where the user is looking, offering a hands-free and intuitive interface for content manipulation.
@@ 3.4 Gaze-Driven Visual Content Generation
Beyond simply recognizing objects users are looking at, we ask: Can we create and edit visual content using just our eyes? GazeGen enables dynamic visual content generation and editing, leveraging gaze as a natural command. This makes the process more efficient and closely aligned with user intentions. GazeGen incorporates both gaze-driven image editing and video generation, utilizing two key diffusion processes: forward diffusion, which transforms the input image into noisy latent layers for preparation, and reverse diffusion, which refines these layers and integrates the modifications. For clarity, latent layers result from forward diffusion, while refinement and integration are done via reverse diffusion.
@@
We introduce gaze-driven operations such as Addition, Deletion/Replacement, Repositioning, and Material Transfer, facilitating intuitive editing of visual content, gaining insights from recent advancements in image editing. By leveraging gaze to obtain object masks and bounding boxes through FastSAM or SAM, and Multimodal Large Language Models (MLLM) respectively, we achieve accurate and user-aligned editing.
Addition. To incorporate new objects based on the user’s gaze, we use MLLM (e.g., LLaVA (Liu et al. 2023)) to suggest the bounding box from the user’s gaze point. A diffusion model then generates the object within this specified area, followed by segmentation (e.g., SAM (Kirillov et al. 2023)) to create a mask that isolates its latent representation. This masked latent representation is subsequently blended with the latent space of the original image, ensuring the new object integrates naturally with the existing visual content.
Deletion/Replacement. For deletion, FastSAM (Zhao et al. 2023) retrieves the object mask within the gaze-specified region. The latent layers corresponding to this region are removed and replaced with Gaussian noise. A reverse diffusion process then regenerates these areas to ensure a coherent image. For replacement, a new object can be added to the removed area using the addition process.
Repositioning. Furthermore, we can perform repositioning by aligning objects with new positions determined by tracking multiple gaze points. To facilitate this, FastSAM retrieves the object mask to define the boundaries. The segmented object is then shifted in the latent space to its new position, while the latent layers at the original position are removed and replaced with Gaussian noise. A reverse diffusion process then regenerates these areas to ensure a coherent background, while also refining the object boundary and seamlessly integrating it into its new location.
Material Transfer. This process begins with using eye gaze to extract an object as a material exemplar, which is encoded for use in a diffusion model. The encoded features are integrated into the diffusion model via cross-attention layers for additional generation conditions. Concurrently, the user’s viewed image is converted into a depth map, and another tracked gaze is used to obtain an object mask. These elements provide geometry and illumination guidance. The depth-based diffusion model offers geometric guidance, while the diffusion model uses these features and the object mask to transfer material properties onto the target object, preserving other attributes.
@@
Addition. To incorporate animated objects into a T2V model using gaze, we utilize a combination of backward and forward diffusion processes. Initially, the user’s viewed image is encoded into spatio-temporal latent codes, $z_{0}^{1:f}$, using a pre-trained T2V model with $f$ repeated frames in the animation. For the region indicated by the bounding box, this area is replaced with Gaussian noise to prepare it for new content.
The diffusion process is represented as:
$z_{t}^{1:f}=\sqrt{\alpha_{t}}z_{0}^{1:f}+\sqrt{1-\alpha_{t}}\epsilon^{1:f},\quad\epsilon\sim\mathcal{N}(0,I),$
where $z_{0}^{1:f}$ denotes the initial spatio-temporal latent codes, $\alpha_{t}$ controls the noise level at time step $t$, and $\epsilon$ represents Gaussian noise.
Subsequently, reverse diffusion is employed to generate animated objects that blend with the spatio-temporal latents. This process ensures that the newly introduced animated content integrates naturally into the existing visual context, resulting in a cohesive and dynamic animation.
Replacement. For replacement, the process begins by retrieving an object mask within the gaze-specified region using FastSAM. The latent layers corresponding to this region are then removed and replaced with Gaussian noise. A reverse diffusion process regenerates these areas to ensure coherent frames. Subsequently, a new animated object is added to the removed area using the addition process described above.
@@ 3.5 GazeGen in Practice
Gaze-Driven Visual Content Generation/Detection. We leverage advanced models to achieve training-free gaze-driven visual content generation and detection, enabling intuitive user interactions. For image editing, objects are added based on bounding boxes suggested by LLaVA (Liu et al. 2023), while FastSAM (Zhao et al. 2023) retrieves object masks for removal. For material transfer, we utilize a diffusion model combined with geometric guidance techniques to ensure accurate transformations. The T2V diffusion model generates animations from gaze-driven inputs, enabling dynamic content creation. Additionally, YOLOv9 (Wang, Yeh, and Liao 2024) identifies and classifies objects within the scene, facilitating gaze-driven object detection.
@@ 4.2 Teacher-Student Model Comparison
Figure 6: Qualitative results for gaze-driven image editing. The tasks include: Addition (first row): Adding objects like a lantern, basket, or photo. Deletion/Replacement (second row): Replacing objects with items like a curtain, aquarium, or galaxy. Reposition (third row): Moving objects such as a wall decoration to the upper left corner, books to the lower left corner, or a phone upward. Material Transfer (last row): Changing an object’s style, such as polished wood to the fridge, woven wicker to the washing machine, or polished metal to the chopping board. All edits are based on the user’s gaze.
@@ 4.4 Qualitative Results
Gaze-Driven Image Editing. Next, we present results from various gaze-driven image editing tasks as shown in Fig. 6. Addition: The first row shows how objects like a lantern, basket, or photo are added to the scene based on where the user looks, enhancing the environment interactively. Deletion/Replacement: In the second row, objects are replaced or removed, such as swapping out items for a curtain, aquarium, or galaxy. This demonstrates the system’s ability to dynamically transform the visual context. Reposition: The third row illustrates repositioning, where objects like a wall decoration are moved to new locations, such as the upper left corner, books to the lower left corner, or a phone moved upward, all guided by the user’s gaze. Material Transfer: The final row demonstrates changing the material style of objects based on the user’s gaze. For instance, the style of the first object seen by the user is applied to the next object they look at. Examples include applying a polished wood texture to a fridge, woven wicker to a washing machine, or polished metal to a chopping board. These changes reflect how gaze can influence the aesthetic and functional attributes of objects.
@@ 7 Conclusion
This paper introduces GazeGen, a hands-free system for visual content generation using eye gaze, enhancing user engagement and accessibility in AR environments. At its core is the DFT Gaze agent, an ultra-lightweight model with 281K parameters, delivering real-time, accurate gaze predictions. It elevates eye gaze from basic tracking to dynamic visual manipulation, enabling tasks like adding, deleting, repositioning elements, material transfer, and converting static images into videos. We developed a compact gaze estimation model using knowledge distillation and a masked autoencoder, refined with Adapters for precise, personalized gaze predictions. These predictions allow GazeGen to facilitate intuitive visual content manipulation and real-time object detection by targeting regions of interest indicated by the user’s gaze, thus enhancing responsiveness and the creative process. Overall, GazeGen sets a new standard for gaze-driven visual content generation, positioning users as active creators and broadening the scope of gaze-driven interfaces.

### v2

@@ Abstract
We present GazeGen, a user interaction system that generates visual content (images and videos) for locations indicated by the user’s eye gaze. GazeGen allows intuitive manipulation of visual content by targeting regions of interest with gaze. Using advanced techniques in object detection and generative AI, GazeGen performs gaze-controlled image adding/deleting, repositioning, and surface style changes of image objects, and converts static images into videos. Central to GazeGen is the DFT Gaze (Distilled and Fine-Tuned Gaze) agent, an ultra-lightweight model with only 281K parameters, performing accurate real-time gaze predictions tailored to individual users’ eyes on small edge devices. GazeGen is the first system to combine visual content generation with real-time gaze estimation, made possible exclusively by DFT Gaze. This real-time gaze estimation enables various visual content generation tasks, all controlled by the user’s gaze. The input for DFT Gaze is the user’s eye images, while the inputs for visual content generation are the user’s view and the predicted gaze point from DFT Gaze. To achieve efficient gaze predictions, we derive the small model from a large model (10x larger) via novel knowledge distillation and personal adaptation techniques. We integrate knowledge distillation with a masked autoencoder, developing a compact yet powerful gaze estimation model. This model is further fine-tuned with Adapters, enabling highly accurate and personalized gaze predictions with minimal user input. DFT Gaze ensures low-latency and precise gaze tracking, supporting a wide range of gaze-driven tasks in AR environments. Leveraging these precise gaze predictions, GazeGen facilitates visual content generation through diffusion processes, allowing users to intuitively manipulate visual content by targeting regions with their gaze. Additionally, it enables real-time object detection by focusing on specific regions indicated by the user’s gaze, improving responsiveness. We validate the performance of DFT Gaze on AEA and OpenEDS2020 benchmarks, demonstrating low angular gaze error and low latency on the edge device (Raspberry Pi 4). Furthermore, we describe applications of GazeGen, illustrating its versatility and effectiveness in various usage scenarios.
@@
Figure 2: Extended applications of gaze-driven interaction with GazeGen. (1) Real-Time Gaze Estimation: Continuous tracking of eye movements for precise gaze estimation. (2) Gaze-Driven Detection: Detecting and identifying objects based on where the user is looking. (3) Gaze-Driven Image Editing: Dynamic editing tasks such as Addition (adding objects based on the user’s gaze), Deletion/Replacement (removing or replacing objects based on the user’s gaze), Reposition (move objects by first gazing at the initial position, then the new position), and Style Transfer (change an object’s style or texture by first gazing at a reference object, then applying the style to the target object). (4) Gaze-Driven Video Generation: Creating and manipulating video content driven by the user’s gaze.
@@ 1 Introduction
With this robust gaze estimation foundation, GazeGen extends its capabilities to real-time object detection by leveraging gaze points to focus on specific regions of the image, retrieving object categories and bounding boxes. For visual content generation, GazeGen uses gaze as a natural command for dynamic image editing and video creation, enabling intuitive operations such as addition, deletion, repositioning, and style transfer. This comprehensive approach allows users to seamlessly manipulate visual content through their gaze, setting a new standard for accessibility and efficiency in the field.
@@
    Use of Eye Gaze for Visual Content Manipulation: We are the first to propose using eye gaze for comprehensive visual content generation and editing, such as adding, deleting, repositioning elements, style transfer, and generating videos. Additionally, GazeGen can detect and interact with objects based on where the user is looking, offering a hands-free and intuitive interface for content manipulation.
@@ 3.4 Gaze-Driven Visual Content Generation
Beyond simply recognizing objects users are looking at, we ask: Can we create and edit visual content using just our eyes? GazeGen enables dynamic visual content generation and editing, leveraging gaze as a natural command. This makes the process more efficient and closely aligned with user intentions. GazeGen incorporates both gaze-driven image editing and video generation, utilizing forward diffusion and reverse diffusion.
@@
We introduce gaze-driven operations such as Addition, Deletion/Replacement, Repositioning, and Style Transfer, facilitating intuitive editing of visual content, gaining insights from recent advancements in image editing.
Addition. To incorporate new objects based on the user’s gaze, we use MLLM (e.g., LLaVA (Liu et al. 2023)) suggests the bounding box from the user’s gaze point, and generative AI synthesizes the object within the specified area.
Deletion/Replacement. For deletion, the object area is removed, and generative AI regenerates the region to ensure a coherent image. For replacement, generative AI synthesizes a new object within the same area.
Repositioning. Repositioning is achieved by tracking multiple gaze points to determine the new position for an object. The object is moved to its new location, and the original area is filled to ensure a consistent background. Generative AI refines the object and blends it into its new surroundings.
Style Transfer. The process uses eye gaze to guide the extraction and transfer of style onto a target object using generative AI.
@@
Addition. To incorporate animated objects into a T2V model using gaze, reverse diffusion is employed to generate cohesive and dynamic animations.
Replacement. For replacement, the object’s area is removed, and generative AI synthesizes an animated object.
@@ 3.5 GazeGen in Practice
Gaze-Driven Visual Content Generation/Detection. We leverage advanced models to achieve training-free gaze-driven visual content generation and detection, enabling intuitive user interactions. For image editing, objects are added based on bounding boxes suggested by LLaVA (Liu et al. 2023). Additionally, YOLOv9 (Wang, Yeh, and Liao 2024) identifies and classifies objects within the scene, facilitating gaze-driven object detection.
@@ 4.2 Teacher-Student Model Comparison
Figure 6: Qualitative results for gaze-driven image editing. The tasks include: Addition (first row): Adding objects like a lantern, basket, or photo. Deletion/Replacement (second row): Replacing objects with items like a curtain, aquarium, or galaxy. Reposition (third row): Moving objects such as a wall decoration to the upper left corner, books to the lower left corner, or a phone upward. Style Transfer (last row): Changing an object’s style, such as polished wood to the fridge, woven wicker to the washing machine, or polished metal to the chopping board. All edits are based on the user’s gaze.
@@ 4.4 Qualitative Results
Gaze-Driven Image Editing. Next, we present results from various gaze-driven image editing tasks as shown in Fig. 6. Addition: The first row shows how objects like a lantern, basket, or photo are added to the scene based on where the user looks, enhancing the environment interactively. Deletion/Replacement: In the second row, objects are replaced or removed, such as swapping out items for a curtain, aquarium, or galaxy. This demonstrates the system’s ability to dynamically transform the visual context. Reposition: The third row illustrates repositioning, where objects like a wall decoration are moved to new locations, such as the upper left corner, books to the lower left corner, or a phone moved upward, all guided by the user’s gaze. Style Transfer: The final row demonstrates changing the style of objects based on the user’s gaze. For instance, the style of the first object seen by the user is applied to the next object they look at. Examples include applying a polished wood texture to a fridge, woven wicker to a washing machine, or polished metal to a chopping board. These changes reflect how gaze can influence the aesthetic and functional attributes of objects.
@@ 7 Conclusion
This paper introduces GazeGen, a hands-free system for visual content generation using eye gaze, enhancing user engagement and accessibility in AR environments. At its core is the DFT Gaze agent, an ultra-lightweight model with 281K parameters, delivering real-time, accurate gaze predictions. It elevates eye gaze from basic tracking to dynamic visual manipulation, enabling tasks like adding, deleting, repositioning elements, style transfer, and converting static images into videos. We developed a compact gaze estimation model using knowledge distillation and a masked autoencoder, refined with Adapters for precise, personalized gaze predictions. These predictions allow GazeGen to facilitate intuitive visual content manipulation and real-time object detection by targeting regions of interest indicated by the user’s gaze, thus enhancing responsiveness and the creative process. Overall, GazeGen sets a new standard for gaze-driven visual content generation, positioning users as active creators and broadening the scope of gaze-driven interfaces.